OpenStudy (anonymous):

Solve: 2(x + 4) − (x + 1) = 10 Anyone wanna help?

OpenStudy (anonymous):

Its 3? I got 2.

OpenStudy (dhernandezcruz):

First, distribute the 2 into the x and 4 and distribute the negative 1 into the x and 1. so you get : 2(x+4) - (x+1)=10 2x+8-x-1= 10 then group like terms: x+7=10 finally subtract 7 from 10 to get : x=3.

OpenStudy (dhernandezcruz):

show me what you did @katrynakay

OpenStudy (anonymous):

 @dhernandezcruz

OpenStudy (dhernandezcruz):

you did't include the -1 when you distibuted the negative 1. you just wrote: 2x+8-1x=10 instead of: 2x+8-1x-1=10.

OpenStudy (dhernandezcruz):

@katrynakay

OpenStudy (anonymous):

OHH. thankyou. lol that makes more sense now.

OpenStudy (dhernandezcruz):

you are very welcome